OpenStudy (anonymous):

√2 (8√t - 5)? Simplify if Possible

OpenStudy (anonymous):

\[\sqrt{2}(8\sqrt{t} - 5)\] Is this the problem?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

If you would like me to type them out, there are multiple choice.

OpenStudy (anonymous):

Just ditribute the √2. \[\sqrt{2}(8\sqrt{t} - 5) = 8\sqrt{t * 2} - 5\sqrt{2} = 8\sqrt{2t} - 5\sqrt{2}\]

OpenStudy (anonymous):

Yeah I did, and I ended up with 8√2t - 5√2; I just wasn't positive if that was its simplest form